# Decoding AI Judgment: How LLMs Assess News Credibility and Bias

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly embedded in workflows that involve evaluative processes. This raises the need to examine how such evaluations are built, what assumptions they rely on, and how their strategies diverge from those of humans. We benchmark six LLMs against expert ratings—NewsGuard and Media Bias/Fact Check (MBFC)—and against human judgments collected through a controlled experiment. To enable direct comparison, we implement a structured agentic framework in which both models and non-expert participants follow the same evaluation procedure: selecting criteria, retrieving content, and producing justifications. Despite output alignment, LLMs rely on different mechanisms: lexical associations and statistical priors replace contextual reasoning. This reliance produces systematic effects: political asymmetries, opaque justifications, and a tendency to confuse linguistic form with epistemic validity. Delegating judgment to such systems does not merely automate evaluation—it redefines it, shifting from normative reasoning to pattern-based approximation.
@@ Introduction
Large Language Models (LLMs) are increasingly embedded in workflows involving classification, evaluation, recommendation, and decision support [1]. Beyond assistive tools, they may influence institutional decision-making, where automated outputs support high-stakes judgments. Advances like chain-of-thought prompting [2] and semi-autonomous agents [3, 4] mark a broader shift: we are no longer just automating tasks, but embedding evaluative functions into socio-technical systems [5]. This raises a key question: what heuristics are encoded when decisions are delegated to LLMs, and how are classifications produced, justified, and interpreted? LLMs can produce outputs similar to those of humans in structured tasks [6], but the similarity concerns results, not the process. We analyze how LLMs apply the concepts of reliability and bias, normative categories that influence which content is shown, hidden, or ignored. These classifications shape information exposure, platform moderation, and public trust. Understanding how models handle these constructs is necessary to assess their societal and epistemic impact. To investigate this, we provide a benchmark involving six LLMs, two expert rating systems, and a sample of human evaluators. Domains are classified by reliability and political bias, allowing us to compare both outputs and the procedures behind them. This helps show how automated evaluations change not just efficiency, but also the criteria and assumptions used in decision-making. LLMs often use surface-level text patterns instead of reasoning from evidence [7, 8]. We compare models and humans not just on their answers, but on how they get there. In online spaces where information spreads fast and users are split into echo chambers [9, 10, 11, 12, 13], judging source credibility is a core problem [14, 15, 16]. These judgments shape what people believe and how public debate evolves—driving polarization, spreading misinformation, and eroding trust [17, 18, 19, 20, 21]. Humans tend to rely on basic standards: is it accurate, independent, and transparent? [22, 23, 24]
These systems—like NewsGuard and MBFC—apply clear rules to rate thousands of news sites [25, 26, 27]. They weren’t built for automation, but they work well as benchmarks to see how LLMs handle reliability and bias.
Our goal is not to evaluate whether LLMs can replace human raters, but to use these benchmarks to examine the heuristics they rely on. Models like GPT [28], Gemini [29], and Llama [30]—already applied in classification and fact-checking tasks [31, 32, 33, 34, 35, 36, 37, 38, 39]—provide a natural testbed for this analysis.
Recent studies have explored whether LLMs replicate human heuristics [40, 3, 41] or reflect ideological biases learned during training [42, 43, 44], but most focus on output-level metrics such as accuracy or bias. Few examine how these outputs are produced.
We address this gap by analyzing how LLMs generate judgments about reliability and political bias, and how their procedures compare to human evaluation.
Six LLMs classify 2,286 domains by trustworthiness and political leaning. Their outputs are compared to expert ratings, and we examine the lexical patterns and explanatory cues they provide to infer the underlying heuristics.
We do not assume human-like reasoning. Instead, we treat model decisions as outcomes of statistical associations shaped by training and prompting. The aim is to understand how LLMs apply normative categories, even when their results align with human judgments.
To extend the analysis, we implement a structured protocol where LLMs simulate evaluative behavior: selecting criteria from a predefined set, retrieving content, and producing justifications. Human participants follow the same procedure in a controlled setting.
While model outputs often align with expert ratings, the underlying processes differ. LLMs rely on surface cues—keywords, frequency patterns, and associative signals—rather than contextual understanding or explicit standards. Human evaluators draw on domain knowledge and normative criteria such as editorial practices and transparency.
In many cases, LLMs generate consistent classifications without accessing content, using only prior associations. This asymmetry shows that delegating evaluations to models changes how reliability and bias are assessed—shifting from human judgment to statistical approximation.
@@
To investigate how LLMs perform in practice, we begin by analyzing their classification of news domains along the dimensions of reliability and political bias. Specifically, we evaluate six state-of-the-art models—Deepseek V3, Gemini 1.5 Flash, GPT-4o mini, Llama 3.1 405B, Llama 4 Maverick, and Mistral Large 2—by comparing their outputs to expert human benchmarks from NewsGuard and MBFC. Beyond assessing alignment with expert judgments, we aim to examine the heuristics and decision patterns these models deploy. Rather than focusing solely on classification accuracy, we investigate the underlying processes that guide model evaluations. To support this analysis, we construct a diverse and balanced dataset of 7,715 English-language news domains. The sample spans multiple countries and includes outlets with both national and international reach.
We extract a snapshot of each domain’s homepage, removing nonessential elements (e.g., scripts, styling) to isolate relevant textual content such as headlines and descriptions. This pre-processing step ensures that all LLMs are evaluated on the same contextual input a human assessor would plausibly consider. The final dataset includes 2,286 active domains successfully classified and output in well-formed JSON documents by all models.
A detailed breakdown of the data collection and processing is provided in the Methods section.
Beyond analyzing the final labels produced by the models, we examine the strategies they use to generate these judgments, thus providing insight into how LLMs encode and apply the notion of reliability.
We begin our assessment by querying each model using a zero-shot, closed-book approach, providing no examples or explicit definitions of reliability. This setup forces models to rely solely on internal representations acquired during training. Our goal is to probe the heuristics LLMs implicitly use when classifying news outlets, and to evaluate how their judgments align with or diverge from structured human evaluations.
To move beyond a simple binary classification (Reliable or Unreliable), we prompt each model to justify its assessment by generating explanatory keywords and assigning a political orientation label to each outlet. Using a standardized prompt across all six models, this setup enables direct comparison of their outputs and allows us to examine how reliability judgments are constructed and where they diverge from human evaluations.
@@
Figure 1: LLMs’ classification against expert human evaluators. (a) Each panel compares how domains rated as “Reliable” or “Unreliable” by NewsGuard are classified by each LLM (Deepseek V3, Gemini 1.5 Flash, GPT-4o mini, Llama 3.1 405B, Llama 4 Maverick, Mistral Large 2). All six models accurately identify “Unreliable” sources, with agreement ranging from $85$ to $97\%$ across models. However, “Reliable” domains show greater classification variability, particularly in Llama 4 Maverick, which classifies a significant portion (35%) as “Unreliable”. (b) We randomly sample 40 domains from each pairing of NewsGuard’s political orientation and reliability rating, and bootstrapped average misclassification rate over 10,000 resamples of reliability misclassification for each. The error bars report the first and third quartile of the resulting frequencies per group. Compared with NewsGuard, LLMs appear to overestimate or underestimate the reliability of news outlets based on their political orientation. In particular, Right-leaning news outlets tend to be consistently misclassified by the LLMs as unreliable, whereas the Center and Left-leaning as reliable.
@@
Figure 1a compares the classifications produced by each model with the reliability ratings assigned by NewsGuard. These ratings are not arbitrary but derived from a structured evaluation protocol based on systematic assessments of editorial standards, transparency, and factual accuracy. By contrast, LLMs make decisions without access to these guidelines, relying instead on internal heuristics formed during training.
@@
All six models accurately identify “Unreliable” sources, consistently flagging domains that NewsGuard associates with low credibility or lack of transparency. In contrast, classifying Reliable sources proves more difficult. GPT-4o mini and Llama 4 Maverick, in particular, misclassify 32% and 35% of reliable domains, respectively—substantially more than the other models. This asymmetry likely reflects the multifaceted nature of NewsGuard’s evaluation criteria (e.g., editorial standards, correction policies, ownership transparency), which may not be fully captured from homepage content alone. We further assess model alignment with expert judgments by comparing their classifications against the Credibility ratings assigned by Media Bias Fact Check (MBFC) for a subset of 923 overlapping domains.
@@
Consistent with earlier results, LLMs show strong agreement with MBFC ratings at the extremes: over 90% of sources labeled Low or High credibility are correctly classified as Unreliable and Reliable, respectively. For Medium credibility sources, however, model performance diverges—both from MBFC and across models. For instance, GPT-4o mini and Llama 4 Maverick classify the majority of these domains as Unreliable (75% and 77%, respectively), while Gemini 1.5 Flash produces a more balanced distribution. This pattern suggests that LLMs rely on clear-cut textual cues when available, but struggle with ambiguous or borderline cases. The confusion matrices for each model are provided in Supplementary Figure S1.
Although the models lack explicit access to the evaluation procedures of NewsGuard and MBFC, and are not provided with their methodological criteria, their outputs reveal consistent lexical associations that approximate the credibility standards applied by human fact-checkers.
@@
As shown in Fig.1b, classification errors are not evenly distributed across the political spectrum. Among domains rated as Reliable by NewsGuard, Right-leaning outlets are consistently misclassified as Unreliable more often than Center or Left-leaning ones, whose reliability tends instead to be overestimated.
@@
Figure 2: Rank-frequency distributions of keywords used by each LLM to describe domains. Each panel presents the most frequently used classification (a) and determinant (b) keywords for Reliable and Unreliable domains. Only the five most common keywords per panel are labeled to enhance readability. The color gradient represents the inferred political orientation of each keyword, ranging from Left-leaning to Right-leaning, based on the political leaning of the domains they are most frequently associated with. Right-leaning keywords appear almost exclusively in descriptions of Unreliable domains, whereas politically neutral or Left-leaning keywords are more characteristic of Reliable domains. All distributions exhibit heavy-tailed behavior, as indicated by their roughly linear shape on a log-log scale, where a small set of highly frequent keywords dominate the descriptions, while the majority appear less frequently. This indicates that LLMs produce consistent markers when explaining their reliability evaluations.
We examine the factors driving LLMs’ reliability judgments by analyzing three sets of keywords generated by each model for every news outlet, alongside their assigned reliability and political orientation labels. These lexical cues offer insight into how models internally represent credibility in the absence of explicit scoring guidelines. Unlike human evaluators, LLMs rely on implicit heuristics—emerging from patterns in their training data—which makes it crucial to uncover the associations that guide their decisions.
@@
To quantify the political leaning of each keyword, we compute the average orientation of the domains in which it appears, using the political labels assigned by the models. These categorical labels are mapped onto a numerical scale from $-1$ (Left) to $1$ (Right), with intermediate values assigned as follows: $-0.5$ (Center-Left), 00 (Center), and $0.5$ (Center-Right).
@@
To analyze keyword usage, we construct rank-frequency distributions separately for each model, keyword type, and reliability label. In these distributions, terms are ordered by frequency of occurrence, with rank 1 assigned to the most frequent. As in natural language corpora, we observe heavy-tailed patterns—few keywords occur frequently, while most appear rarely—reflecting the typical statistical structure of language.
@@
To streamline the presentation, we focus on three representative models—Gemini, GPT, and Llama 4—with results for the remaining models reported in the Supplementary Information. Figure 2 shows the rank-frequency distributions of “classification” and “determinant” keywords by model and reliability label. All models exhibit a heavy-tailed pattern, indicating reliance on a core set of linguistic markers when evaluating reliability. Further, it suggests that LLMs do not apply a coherent evaluative model, but rely on statistical salience, conflating stylistic markers with epistemic features.
@@
As shown in Fig. 2, classification keywords capture key linguistic markers used by models in reliability assessment. Reliable domains are frequently associated with terms denoting neutrality, transparency, and factual reporting, suggesting a focus on balanced communication and professional presentation. In contrast, unreliable domains are consistently linked to terms such as “misinformation”, “conspiracy”, and “bias”, reflecting associations with sensationalism and partisanship. These patterns suggest that LLMs develop structured linguistic heuristics that partially mirror human evaluative criteria.
@@
Determinant keywords offer further insight into the mechanisms driving model classifications. Reliable domains are often associated with references to editorial standards and institutional transparency. GPT-4o mini and Llama 4 notably emphasize “local news,” suggesting that community-based reporting is interpreted as a marker of credibility. In contrast, unreliable domains are consistently linked to politically charged terms, with keywords such as “trump”, “biden”, “deep state”, and “fake news” recurring prominently—indicating that politicized content is a strong signal for unreliability.
@@
Additionally, Fig. 2 reveals an asymmetry in the political connotation of keywords: right-leaning terms are more prevalent in descriptions of unreliable sources, whereas neutral or left-leaning terms more often appear in association with reliable domains.
@@
Figure 3: Keywords’ rank among “reliable” and “unreliable” domains. We label only keywords sufficiently distant from the diagonal, meaning they are predominantly used to describe reliable or unreliable domains rather than being evenly distributed across both classifications. Additionally, we label the top 5 keywords per reliability rating. The color gradient represents the inferred political orientation of each keyword, from Left-leaning to Right-leaning, based on the domains with which they are most frequently associated. While summary keywords (bottom row) appear with similar frequency in both reliable and unreliable domains, classification and determinant keywords (top and middle rows) exhibit sharper separation. This result suggests that reliable and unreliable sources may cover similar topics but differ in framing tone or contextual emphasis. Notably, keywords related to transparency, objectivity, and credibility are more common among reliable domains. At the same time, sensationalist and politicized terms such as “misinformation”, “propaganda”, and “bias” are frequently linked to unreliable sources.
@@
When examining “classification” and “determinant” keywords, clear differences emerge between the two reliability groups. Reliable domains are associated with terms such as “local news”, “scientific”, “diverse”, and “data-driven”, while unreliable classifications involve more controversial or politically charged terms, including politician names (e.g., “trump”, “biden”) and topics such as “genocide” and ”vaccines”. In contrast, “summary” keywords—describing the overall content of a domain—show substantial overlap between the two groups. This suggests that reliable and unreliable sources often cover similar topics, and that the distinction lies less in what is covered than in how it is presented. We also find that terms with no clear semantic polarity tend to be unevenly distributed across classes, revealing that model judgments are influenced by framing and usage context rather than meaning alone.
These findings suggest that LLMs do not assess news outlets by applying explicit criteria. Instead, they rely on implicit heuristics shaped by training data, drawing on linguistic framing, stylistic regularities, and contextual cues rather than explicit factual content. This raises important questions about how credibility is operationalized in these models and whether their evaluative strategies reflect patterns learned from human discourse.
@@
Figure 4: Reliability evaluations by Gemini-powered LLM agents and non-expert humans in a controlled experimental setting. (a) The two panels compare humans’ and agents’ reliability ratings against NewsGuard’s classifications. Models consistently identify all Unreliable (U) sources and struggle with the Reliable (R). In contrast, humans show little to no alignment with NewsGuard, for both reliable and unreliable domains. (b) Confusion matrix of ratings provided by humans and agents, with the human ratings used as the ground truth. The two show strong agreement on unreliable sources, while $80\%$ of sources rated as reliable by humans are considered unreliable by the LLM. (c) Distributions of order choices for each criterion by humans (left) and models (right). The human distributions appear more uniform than those of the models, indicating that each criterion is roughly equally likely to appear in any position compared to LLMs.
Our previous analysis shows that LLMs often produce reliability ratings that align closely with expert evaluations from NewsGuard and MBFC. This suggests that, despite lacking access to structured evaluation criteria, these models have internalized heuristics that approximate human judgments. Yet a key question remains: what cognitive procedures or proxies underlie these evaluations?
@@
This raises a foundational question: are LLMs performing actual input-dependent evaluation or simply retrieving associations learned during training? If a model can classify a domain without analyzing its content, it becomes difficult to disentangle reasoning from statistical recall. More broadly, such behavior suggests that these models may not be applying evaluative criteria in a structured or integrative fashion, but rather relying on fragmentary heuristics shaped by training data.
@@
To probe this issue, we introduce a structured agentic workflow that enables a direct comparison between LLMs and human evaluators. Rather than treating models as black boxes producing binary outputs, we instantiate a multi-agent system designed to simulate the procedural steps involved in human evaluation—retrieving, processing, and integrating information before rendering a judgment. This design allows us to assess whether LLMs can apply criteria in a coordinated sequence, or whether they continue to operate in a cue-based mode. For this analysis, we focus on Gemini 2.0 Flash, which optimally supports the tooling required for implementing the agentic pipeline. Given the consistency observed across models in prior sections, we expect the insights to generalize.

### v3

# The simulation of judgment in LLMs

@@ Abstract
Please refer to published version: https://doi.org/10.1073/pnas.2518443122
Large Language Models (LLMs) are increasingly embedded in evaluative processes, from information filtering to assessing and addressing knowledge gaps through explanation and credibility judgments. This raises the need to examine how such evaluations are built, what assumptions they rely on, and how their strategies diverge from those of humans. We benchmark six LLMs against expert ratings—NewsGuard and Media Bias/Fact Check—and against human judgments collected through a controlled experiment. We use news domains purely as a controlled benchmark for evaluative tasks, focusing on the underlying mechanisms rather than on news classification per se. To enable direct comparison, we implement a structured agentic framework in which both models and nonexpert participants follow the same evaluation procedure: selecting criteria, retrieving content, and producing justifications. Despite output alignment, our findings show consistent differences in the observable criteria guiding model evaluations, suggesting that lexical associations and statistical priors could influence evaluations in ways that differ from contextual reasoning. This reliance is associated with systematic effects: political asymmetries and a tendency to confuse linguistic form with epistemic reliability—a dynamic we term epistemia, the illusion of knowledge that emerges when surface plausibility replaces verification. Indeed, delegating judgment to such systems may affect the heuristics underlying evaluative processes, suggesting a shift from normative reasoning toward pattern-based approximation and raising open questions about the role of LLMs in evaluative processes.
@@ Introduction
Large Language Models (LLMs) are increasingly embedded in workflows involving classification, evaluation, recommendation, and decision support [1]. This rapid integration of AI technologies is not just reshaping industries but also presenting a fundamental choice between a path of pure automation and one of human complementation, where technology is designed to augment human spread and mitigate widespread unemployment [2]. Beyond assistive tools, they may influence institutional decision-making, where automated outputs support high-stakes judgments. Advances like chain-of-thought prompting [3] and semi-autonomous agents [4, 5] mark a broader shift: we are no longer just automating tasks, but embedding evaluative functions into socio-technical systems [6]. This delegation carries significant risks, as AI systems trained on biased historical data may replicate and amplify societal inequalities in critical domains, such as the labor market [7].
As these systems scale, the issue is no longer only whether outputs are correct, but how the very notion of judgment is operationalized once decisions are delegated to statistical models. This raises a key question: what heuristics are encoded when decisions are delegated to LLMs, and how are classifications produced, justified, and interpreted? LLMs can produce outputs similar to those of humans in structured tasks [8], but the similarity concerns results, not the process. What appears as alignment at the output level may conceal a deeper epistemic shift, where normative reasoning is replaced by surface-level approximation.
@@
We analyze how LLMs apply the concepts of reliability and bias, normative categories that influence which content is shown, hidden, or ignored. These classifications shape information exposure, platform moderation, and public trust. Understanding how models handle these constructs is necessary to assess their societal and epistemic impact. To investigate this, we provide a benchmark involving six LLMs, two expert rating systems, and a sample of human evaluators. News outlets are classified by reliability and political bias, allowing us to compare both outputs and the procedures behind them. The news domain serves here only as a controlled testbed, enabling us to isolate and analyze the mechanisms behind automated evaluations rather than focusing on domain-specific accuracy. This helps show how automated evaluations change not just efficiency, but also the criteria and assumptions used in decision-making. Indeed, LLMs often use surface-level text patterns instead of reasoning from evidence [9, 10].
In online spaces where information spreads fast and users are split into echo chambers [11, 12, 13, 14, 15], judging source credibility as well as their biases is a core problem [16, 17, 18]. These judgments shape what people believe and how public debate evolves—driving polarization, spreading misinformation, and eroding trust [19, 20, 21, 22, 23]. Humans tend to rely on basic standards: is it accurate, independent, and transparent? [24, 25, 26] Expert human evaluators—like NewsGuard and MBFC—apply clear rules to rate thousands of news sites [27, 28, 29]. Our goal is not to assess whether LLMs can replace human raters, but to use these benchmarks to analyze the heuristics guiding their evaluations and how they operationalize concepts such as credibility and bias. Widely used models such as GPT [30], Gemini [31], and Llama [32]—already employed in numerous classification and fact-checking tasks [33, 34, 35, 36, 37, 38, 39, 40, 41]—offer a natural testbed for this analysis. Recent studies have explored whether LLMs replicate human heuristics [42, 4, 43] or reflect ideological biases learned during training [44, 45, 46], but most focus on output-level metrics such as accuracy or bias. Few examine how these outputs are produced. We address this gap by analyzing how LLMs generate judgments about reliability and political bias, and how their procedures compare to human evaluation. Six LLMs classify 2,286 domains by reliability and political leaning. Their outputs are compared to expert ratings, and we examine the lexical patterns and explanatory cues they provide to infer the underlying heuristics. We do not assume human-like reasoning. Rather, we empirically analyze how models operationalize evaluative tasks, allowing us to test whether their decisions rely on statistical associations shaped by training and prompting. This leads us to ask whether delegating judgment to LLMs preserves its normative and epistemic meaning, or whether it transforms it into what we call epistemia—a condition in which the appearance of coherent and authoritative judgment arises from statistical patterning alone, producing the illusion of knowledge when surface plausibility substitutes for evidence-based reasoning.
Our findings show that model outputs often align with expert ratings of reliability and bias, yet systematic asymmetries emerge across the political spectrum. Moreover, LLMs generate consistent linguistic markers when explaining their evaluations. To test these differences, we implement a structured protocol in which LLMs simulate evaluative behavior—selecting criteria from a predefined set, retrieving content, and producing justifications—while human participants follow the same procedure in a controlled setting. The results show that LLMs and humans prioritize different reliability criteria, consistent with a shift from context-dependent, normative reasoning—understood here as the application of explicit quality standards and contextual reasoning rather than implying perfectly rational agents—toward pattern-based approximation.
Overall, our findings indicate that delegating evaluations to models transforms how reliability and bias are assessed, replacing human judgment with statistical approximation.
@@
To investigate how LLMs perform in practice, we begin by analyzing their classification of online news outlets along the dimensions of reliability and political bias. Specifically, we evaluate six state-of-the-art models—Deepseek V3, Gemini 1.5 Flash, GPT-4o mini, Llama 3.1 405B, Llama 4 Maverick, and Mistral Large 2—by comparing their outputs to expert human benchmarks from NewsGuard and MBFC. Beyond assessing alignment with expert judgments, we aim to examine the heuristics and decision patterns these models deploy, providing a first approach to investigate the underlying processes that guide model evaluations. To support this analysis, we construct a diverse dataset of 7,715 English-language news outlet domains. The sample spans multiple countries and includes outlets with both national and international reach.
We extract a snapshot of each domain’s homepage, removing nonessential elements (e.g., scripts, styling) to isolate relevant textual content such as headlines and descriptions. This pre-processing step ensures that all LLMs are evaluated on the same textual input a human assessor would plausibly consider. The final dataset includes 2,286 active domains successfully classified and output in well-formed JSON documents by all models. A detailed breakdown of the data collection and processing is provided in the Methods section.
Beyond analyzing the final labels produced by the models, we examine the strategies they use to generate these judgments, thus providing insight into how LLMs encode and operationalize the notion of reliability.
We begin our assessment by querying each model using a zero-shot, closed-book approach, providing no examples or explicit definitions of reliability. This setup constrains models to rely on internal representations acquired during training. Our goal is to analyze, through model outputs as a proxy, the heuristics guiding their classifications and to evaluate where their judgments align with or diverge from structured human evaluations.
To move beyond a simple binary classification (Reliable or Unreliable), we prompt each model to assign a political orientation label to each outlet and justify its assessment by generating explanatory keywords. Using a standardized prompt across all six models, this setup enables direct comparison of their outputs and provides insight into how models construct their reliability judgments and how these compare to expert human evaluations.
@@
Figure 1: LLMs’ classification against expert human evaluators. (a) Each panel compares how domains rated as “Reliable” or “Unreliable” by NewsGuard are classified by each LLM (Deepseek V3, Gemini 1.5 Flash, GPT-4o mini, Llama 3.1 405B, Llama 4 Maverick, Mistral Large 2). All six models accurately identify Unreliable sources, with agreement ranging from 85 to 97% across models. However, Reliable domains show greater classification variability, particularly in Llama 4 Maverick and in GPT 4o-mini, which classify a significant portion (35% and 32%) as “Unreliable.” (b) We randomly sample 40 domains from each pairing of NewsGuard’s political orientation and reliability rating and compute the average misclassification rate across political orientations over 10,000 resamples. The error bars report the first and third quartiles of the resulting frequencies per group. Compared with NewsGuard, LLMs appear to overestimate or underestimate the reliability of news outlets based on their political orientation. In particular, Right-leaning news outlets tend to be consistently misclassified by the LLMs as unreliable, whereas the Center and Left-leaning as reliable.
@@
Figure 1a compares the classifications produced by each model with the reliability ratings assigned by NewsGuard. These ratings are not arbitrary but derived from a structured evaluation protocol based on systematic assessments of editorial standards, transparency, and factual accuracy. By contrast, LLMs make decisions without directly accessing these guidelines, relying instead on internal heuristics formed during training.
@@
All six models accurately identify “Unreliable” sources, consistently flagging domains that NewsGuard associates with low credibility or lack of transparency. In contrast, classifying “Reliable” sources proves more difficult. GPT-4o mini and Llama 4 Maverick, in particular, misclassify 32% and 35% of reliable domains, respectively—substantially more than the other models. This asymmetry may reflect the multifaceted nature of NewsGuard’s evaluation criteria (e.g., editorial standards, correction policies, ownership transparency), which may not be fully captured from homepage content alone. We further assess model alignment with expert judgments by comparing their classifications against the Credibility ratings assigned by Media Bias Fact Check (MBFC) for a subset of 916 overlapping domains.
@@
Consistent with Fig. 1a, LLMs often achieve over 90% accuracy with MBFC ratings at the extremes: sources labeled Low or High credibility are correctly classified as Unreliable and Reliable, respectively. For Medium credibility sources, however, model performance diverges—both from MBFC and across models. For instance, GPT-4o mini and Llama 4 Maverick classify the majority of these domains as Unreliable (75% and 77%, respectively), while Gemini 1.5 Flash produces a more balanced distribution. This pattern is consistent with the interpretation that LLMs rely primarily on clear-cut textual cues when available, while showing lower accuracy on ambiguous or borderline cases. The confusion matrices for each model are provided in Supplementary Figure S1. Although the models lack explicit access to the evaluation procedures of NewsGuard and MBFC, and are not provided with their methodological criteria, their outputs are broadly consistent with the credibility assessments made by expert human fact-checkers.
@@
As shown in Fig.1b, classification errors are not evenly distributed across the political spectrum. Among domains rated as Reliable by NewsGuard, Right-leaning outlets are consistently misclassified as Unreliable more often than Center or Left-leaning ones, whose reliability tends instead to be overestimated. Importantly, this asymmetry does not indicate that LLMs hold partisan preferences. Rather, consistent with recent work on value alignment and political bias in LLMs (e.g., [43, 41]), it likely reflects correlations in the training data—for instance, the co-occurrence of extremist rhetoric and misinformation—rather than an explicit ideological stance. As a result, models may overgeneralize, conflating legitimate right-leaning journalism with toxic or conspiratorial sources when linguistic markers overlap.
@@
Figure 2: Rank-frequency distributions of keywords used by each LLM to describe domains. Each panel presents the most frequently used classification (a) and determinant (b) keywords for Reliable and Unreliable domains. Only the five most common keywords per panel are labeled to enhance readability. The color gradient represents the inferred political orientation of each keyword, ranging from Left-leaning to Right-leaning, based on the political leaning of the domains they are most frequently associated with. Right-leaning keywords appear almost exclusively in descriptions of Unreliable domains, whereas politically neutral or Left-leaning keywords are more characteristic of Reliable domains. All distributions exhibit heavy-tailed behavior, as indicated by their roughly linear shape on a log–log scale, where a small set of highly frequent keywords dominate the descriptions, while the majority appear less frequently. This indicates that LLMs produce consistent markers when explaining their reliability evaluations.
We examine the factors driving LLMs’ reliability judgments by analyzing three sets of keywords generated by each model for every news outlet, alongside their assigned reliability and political orientation labels. These lexical cues provide indirect evidence about the heuristics models used to approximate credibility in the absence of explicit scoring guidelines. Unlike human evaluators, LLMs rely on implicit heuristics—emerging from patterns in their training data—underscoring the importance of analyzing the associations reflected in their outputs.
@@
To quantify the political leaning of each keyword, we compute the average orientation of the domains in which it appears, using the political labels assigned by the models. These categorical labels are mapped onto a numerical scale from 1-1 (Left) to $1$ (Right), with intermediate values assigned as follows: 0.5-0.5 (Center-Left), 0 (Center), and $0.5$ (Center-Right).
@@
To analyze keyword usage, we construct rank-frequency distributions separately for each model, keyword type, and reliability label. In these distributions, terms are ordered by frequency of occurrence, with rank 1 assigned to the most frequent.
@@
To streamline the presentation, we focus on three representative models—Gemini, GPT, and Llama 4—with results for the remaining models reported in the Supplementary Information. Figure 2 shows the rank-frequency distributions of classification and determinant keywords by model and reliability label. All models exhibit a heavy-tailed pattern, indicating that a common core set of linguistic markers is consistently associated with the classification procedure. This is consistent with natural language corpora, where few words occur frequently while most appear rarely [47].
@@
As shown in Fig. 2, classification keywords capture linguistic markers associated with model classifications in reliability assessment. Reliable domains are frequently linked to terms denoting neutrality, transparency, and factual reporting, suggesting a focus on balanced communication and professional presentation. In contrast, unreliable domains are consistently associated with terms such as “misinformation”, “conspiracy”, and “bias”, consistent with patterns often linked to sensationalism and partisanship. These patterns indicate that LLM classifications exhibit structured linguistic heuristics that partially mirror human evaluative criteria.
@@
Determinant keywords offer further insight into the patterns underlying model classifications. Reliable domains are often associated with references to editorial standards and institutional transparency. GPT-4o mini and Llama 4 notably emphasize “local news”, suggesting that community-based reporting appears as a marker of credibility. In contrast, unreliable domains are consistently linked to politically charged terms, with keywords such as “trump”, “biden” and “deep state” recurring prominently, suggesting that politicized content is often associated with unreliability.
@@
Additionally, Fig. 2 highlights an asymmetry in the political connotation of keywords: right-leaning terms are more prevalent in descriptions of unreliable sources, whereas neutral or left-leaning terms appear more often in association with reliable domains.
@@
Figure 3: Keywords’ rank among Reliable and Unreliable domains. We label only keywords sufficiently distant from the diagonal, meaning they are predominantly used to describe reliable or unreliable domains rather than being evenly distributed across both classifications. Additionally, we label the top 5 keywords per reliability rating. The color gradient represents the inferred political orientation of each keyword, from Left-leaning to Right-leaning, based on the domains with which they are most frequently associated. While summary keywords (Bottom row) appear with similar frequency in both reliable and unreliable domains, classification and determinant keywords (Top and Middle rows) exhibit sharper separation. This result suggests that reliable and unreliable sources may cover similar topics but differ in framing tone or contextual emphasis. Notably, keywords related to transparency, objectivity, and credibility are more common among reliable domains. At the same time, sensationalist and politicized terms such as “misinformation,” “propaganda,” and “bias” are frequently linked to unreliable sources.
@@
When examining classification and determinant keywords, clear differences emerge between the two reliability groups. Reliable domains are associated with terms such as “local news”, “scientific”, “diverse”, and “evidence-based”, while unreliable classifications involve more controversial or politically charged terms, including politician names (e.g., “trump”, “biden”) and topics such as “genocide” and ”vaccines”. In contrast, summary keywords—describing the overall content of a domain—show substantial overlap between the two groups. This suggests that reliable and unreliable sources often cover similar topics, and that the distinction lies less in what is covered than in how it is presented. We also find that terms with no clear semantic polarity tend to be unevenly distributed across classes, suggesting that model judgments may be associated with framing and usage context rather than meaning alone.
@@
Figure 4: Reliability evaluations by Gemini-powered LLM agents and non-expert humans in a controlled experimental setting. (a) The two panels compare humans’ and agents’ reliability ratings against NewsGuard’s classifications. Models consistently identify all Unreliable (U) sources and struggle with the Reliable (R). In contrast, humans show little to no alignment with NewsGuard, for both reliable and unreliable domains. (b) Confusion matrix of ratings provided by humans and agents, with the human ratings used as the ground truth. The two show strong agreement on unreliable sources, while 77% of sources rated as reliable by humans are considered unreliable by the LLM. (C) Distributions of order choices for each criterion by humans (Left) and models (Right). The human distributions appear more uniform than those of the models, indicating that most criteria are roughly equally likely to appear in any position compared to LLMs.
Our previous analysis shows that LLMs often achieve high agreement with expert evaluations from NewsGuard and MBFC. This suggests that, despite lacking access to structured evaluation criteria, their outputs reflect heuristics that approximate human judgments. Yet a key question remains: what procedures or proxies underlie these evaluations?
@@
This raises a foundational question: are LLMs engaging in content-specific evaluation or relying primarily on statistical associations learned during training? If a model can classify a domain without analyzing its content, it becomes difficult to disentangle content-based assessment from statistical recall. More broadly, such behavior is consistent with the interpretation that model assessments may be shaped by prior knowledge about the news outlet rather than by content-specific evaluation.
@@
To analyze this issue, we introduce a structured agentic workflow that enables a direct comparison between LLMs and human evaluators. Rather than treating models as black boxes producing binary outputs, we instantiate a multi-agent system designed to simulate the procedural steps involved in human evaluation—retrieving, processing, and integrating information before rendering a judgment. An agentic pipeline is well-suited to this task, as it provides LLMs with access to modular tools (i.e., external functions implemented in code) and enables the composition of deterministic multi-step workflows, where intermediate outputs from one LLM can be passed to the next. To implement this workflow effectively, we focus on Gemini 2.0 Flash, which optimally supports the tooling required for implementing the agentic pipeline. Given the consistency observed across models in prior sections, we expect the insights to generalize.